# Operational research approaches and mathematical models for kidney exchange: A literature survey and empirical evaluation

### v1

@@ 3.2 Introduction to ILP Formulations for KE-Opt
As chains were only introduced in practice later, the version of KE-Opt with chains has received less attention. To address this gap, we also show in Sections 3.3-3.7 how each cycle model can be adapted to a chain model that models a set of cardinality-constrained vertex-disjoint chains, instead. The resulting models are denoted as CF-CHAIN, HCF-CHAIN, EF-CHAIN, EEF-CHAIN and PIEF-CHAIN. For EF-CHAIN and EEF-CHAIN we consider two variants each, which are distinguished by adding suffix -EXP or -MTZ. Throughout the following sections, we assume that $L$ is finite, and we discuss in Appendix (A.7) how each of these models can be adapted to the case where $L=\infty$ (meaning that the maximum chain length is unbounded).
Even though there are no KEPs in practice that allow only for chains, separately studying cycle models and chain models enables a comprehensive view of modelling cycles and chains by merging any cycle model with any chain model to obtain a combined model that incorporates a set of cardinality-constrained vertex-disjoint cycles and chains. This works as follows: simply keep the variables and constraints of both the chosen cycle model and the chain model, sum the objective functions, and for each RDP $v\in\mathcal{R}$, merge the constraint in the cycle model stating that $v$ can be in at most one cycle with the corresponding constraint in the chain model stating that $v$ can be in at most one chain to obtain a constraint ensuring that $v$ can be in at most one cycle or chain.
@@
Type Model Section(s) Literature
Cycles CF-CYCLE 3.3 Abraham et al., (2007) / Roth et al., (2007)
HCF-CYCLE 3.4 Delorme et al., (2023)
EF-CYCLE 3.5 Abraham et al., (2007) / Roth et al., (2007)
EEF-CYCLE 3.6 Constantino et al., (2013)
PIEF-CYCLE 3.7 Dickerson et al., (2016)
Chains CF-CHAIN 3.3 Constantino et al., (2013)
HCF-CHAIN* 3.4 Delorme et al., (2023)
EF-CHAIN-EXP* 3.5 Constantino et al., (2013) & Anderson et al., (2015)
@@
EEF-CHAIN-EXP 3.6/A.1 Anderson et al., (2015)
EEF-CHAIN-MTZ* 3.6/A.1 Mak-Hau, (2017)
PIEF-CHAIN 3.7 Dickerson et al., (2016)
Hybrid EF-HYBRID* 3.5/A.3 Constantino et al., (2013) & Anderson et al., (2015)
@@ 3.3 Cycle Formulation
Let $\mathcal{C}_{\leq K}$ be the set of all cycles $c$ in $\mathcal{G}$ of length at most $K$. For every cycle $c\in\mathcal{C}_{\leq K}$ we assume that the first vertex is the lowest-indexed one (to avoid repetitions due to symmetry). Moreover, let $\mathcal{V}(c)$ and $\mathcal{A}(c)$ denote the sets of vertices and arcs in $c$, and let $\omega_{c}=\sum_{(u,v)\in\mathcal{A}(c)}w_{uv}$ denote the total weight of $c$. We introduce a binary decision variable $x_{c}$ for every $c\in\mathcal{C}_{\leq K}$, taking value $1$ if cycle $c$ is selected, and value 00 otherwise. CF-CYCLE can then be defined as follows:
[…]
@@
The objective function (1) maximises the total weight and constraints (2) enforce that all RDPs are involved in at most one cycle.
Example 1.
(continued) Consider the instance presented in Figure 3 for $K\in\{2,3,4\}$. Set $\mathcal{C}_{\leq K}$ only consists of cycle $\left\langle 1,4,1\right\rangle$ if $K=2$, also contains cycle $\left\langle 2,3,4,2\right\rangle$ if $K=3$, and additionally also cycle $\left\langle 1,2,3,4,1\right\rangle$ if $K=4$. That is, when $K=4$, CF-CYCLE contains $3$ variables, one for each cycle of length at most $4$.
The adaptation of CF-CYCLE to model chains instead of cycles is straightforward, and was first described in a general setting by Constantino et al., (2013). Defining $\mathcal{C}^{\prime}_{\leq L}$ as the set of all chains $c$ in $\mathcal{G}$ of length at most $L$, and defining a binary variable $y_{c}$ for every $c\in\mathcal{C}^{\prime}_{\leq L}$, taking value $1$ if chain $c$ is selected and value 00 otherwise, the following model is obtained:
[…]
@@
In Appendix A.6 we explain that the variables associated to chains of length $1$ can be omitted.
Example 1.
(continued) Consider the instance presented in Figure 3 for $L=3$. We have $\mathcal{C}^{\prime}_{\leq L}=\{\left\langle 5,\tau\right\rangle,\left\langle 6,\tau\right\rangle,\left\langle 5,1,\tau\right\rangle,\left\langle 6,2,\tau\right\rangle,\left\langle 5,1,2,\tau\right\rangle,\left\langle 5,1,4,\tau\right\rangle,\left\langle 6,2,3,\tau\right\rangle\}$. Therefore, CF-CHAIN contains $7$ variables, one for each chain of length at most $3$.
@@
The simplest approach is to enumerate all cycles and/or chains, as did, for instance, Roth et al., (2007), Manlove and O’Malley, (2014), Constantino et al., (2013), Anderson et al., (2015) and Dickerson et al., (2016). However, this approach is only viable when $K$ and $L$ are sufficiently small and the compatibility graph is sufficiently small and sparse. For example, Constantino et al., (2013) showed that CF-CYCLE could solve some low-density instances with $|\mathcal{R}|=1000$, $K=3$ and $L=0$, whereas the number of cycles was too high for high-density instances with $|\mathcal{R}|=50$, $K\geq 4$ and $L=0$. These papers deal with chains in different ways. Roth et al., (2007) considered the cycles-only case. Manlove and O’Malley, (2014) studied the case where $K=L$, and applied the chain-to-cycle transformation, after which CF-CYCLE was used. Constantino et al., (2013) extended this to the case where $K\neq L$, which essentially comes down to using the combined model CF-CYCLE + CF-CHAIN. Finally, Anderson et al., (2015) and Dickerson et al., (2016) combined CF-CYCLE with a different chain model, namely a variant of EF-CHAIN-EXP (see Section 3.5) and PIEF-CHAIN (see Section 3.7), respectively. This was motivated by the fact that the number of variables (i.e., chains) in CF-CHAIN typically dominates the number of variables (i.e., cycles) in CF-CYCLE.
@@
Several other improvements for B&P algorithms were pursued. First, different branching rules have been proposed. For example, Abraham et al., (2007) branched directly on fractional cycle/chain variables, whereas Glorie et al., 2014b branched on the underlying fractional arcs, which ensures that the depth of the branching tree is polynomially bounded. Moreover, Klimentova et al., (2014) introduced a new decomposition model for the master problem, called the Disagreggated Cycle Decomposition Model, which was improved by Riascos-Álvarez et al., (2024) using feedback vertex sets. Furthermore, Lam and Mak-Hau, (2020) and Petris et al., (2024) considered branch-and-price-and-cut algorithms, in which one progressively adds cutting planes to strengthen the LP relaxation of the restricted master problem. Whereas Lam and Mak-Hau, (2020) report that their considered valid inequalities are rather ineffective in strengthening the linear relaxation, the inequalities introduced by Petris et al., (2024) are shown to be more effective. Moreover, Riascos-Álvarez et al., (2024) introduced a new upper bound based on Lagrangian relaxation. Very recently, Arslan et al., (2024) proposed two B&P algorithms (one based on CF-CYCLE+CF-CHAIN and one based on CF-CYCLE+PIEF-CHAIN) that combine many of the improvements that were proposed over the years. As shown also by our computational experiments (see Section 4), these algorithms are the current state-of-the-art for KE-Opt on unweighted instances among all publicly available methods.
Delorme et al., (2023) and Delorme et al., (2024) were the first to apply RCVF to KE-Opt. Less involved than B&P, RCVF is a general technique that can be applied to any ILP model, which works especially well when the model has a tight LP relaxation (as is the case for CF-CYCLE and CF-CHAIN). The idea is to deactivate variables for which, based on their reduced cost, it can be deduced that their value will never be $1$ or higher in a solution whose value is equal to some upper bound. This upper bound is initially set to the value of the LP relaxation rounded down, and then progressively decreased within a destructive bound framework until an optimal solution is found.
@@ 3.4 Half-Cycle Formulation
Formally, a half-cycle $h$ of length $k$, denoted as $h=\left\langle v_{1},v_{2},\ldots,v_{k+1}\right\rangle$, is a subgraph of $\mathcal{G}$ containing $k+1$ RDPs $v_{1},\ldots,v_{k+1}\in\mathcal{R}$ such that there is an arc $(v_{i},v_{i+1})\in\mathcal{A}_{\mathcal{R}}$ for every $i=1,\ldots,k$. Let $v^{s}(h)=v_{1}$ and $v^{e}(h)=v_{k+1}$ denote the start and end vertex of half-cycle $h$, respectively, and let $\mathcal{V}^{m}(h)=\{v_{2},\ldots,v_{k}\}$ denote the set of intermediate vertices (when $k\geq 2$). Moreover, letting $\mathcal{A}(h)$ denote the set of arcs in half-cycle $h$, its total weight is given by $\omega_{h}=\sum_{(i,j)\in\mathcal{A}(h)}w_{ij}$. Using this notation, we have that two half-cycles, say $h_{1}$ of length $k_{1}$ and $h_{2}$ of length $k_{2}$, are compatible (i.e., they may be combined to a full cycle of weight $\omega_{h_{1}}+\omega_{h_{2}}$) if $v^{e}(h_{1})=v^{s}(h_{2})$, $v^{e}(h_{2})=v^{s}(h_{1})$, $\mathcal{V}^{m}(h_{1})\cap\mathcal{V}^{m}(h_{2})=\emptyset$, and $k_{1}+k_{2}\leq K$.
Let $\mathcal{H}$ denote the set of all half-cycles. Given an ordering of the vertices in $\mathcal{V}$, it is sufficient to only consider half-cycles $h$ of length $k$ that satisfy the following conditions: (i) $k\leq\left\lceil K/2\right\rceil$, (ii) either $v^{s}(h)$ or $v^{e}(h)$ is the smallest indexed vertex among vertices in $h$, but it must be $v^{s}(h)$ if $K$ is odd and $k=\left\lceil K/2\right\rceil$, and (iii) there exists at least one compatible half-cycle of length $k$ or $k-1$ if $v^{s}(h)<v^{e}(h)$, or length $k$ or $k+1$ if $v^{s}(h)>v^{e}(h)$ [^3] In our implementation, we first construct all half-cycles satisfying conditions (i) and (ii), after which we remove all half-cycles that do not satisfy a relaxed version of condition (iii). Namely, we ignore the requirement that two compatible half-cycles $h_{1}$ and $h_{2}$ must satisfy $\mathcal{V}^{m}(h_{1})\cap\mathcal{V}^{m}(h_{2})=\emptyset$.. There are multiple ways to determine an ordering of the vertices, but Delorme et al., (2023) proposed to sort the vertices in descending order of total vertex degree, as this typically leads to a smaller number of half-cycles.
Introducing a binary decision variable $x_{h}$ for every half-cycle $h\in\mathcal{H}$, taking value $1$ if half-cycle $h$ is selected, and value 00 otherwise, HCF-CYCLE can be defined as the following ILP:
[…]
s.t. $\sum_{h\in\mathcal{H}:v\in\mathcal{V}^{m}(h)\cup\{v^{s}(h)\}}x_{h}\leq 1$ $\forall v\in\mathcal{R},$ (8)
$\sum_{h\in\mathcal{H}:v^{s}(h)=u,v^{e}(h)=v}x_{h}=\sum_{h\in\mathcal{H}:v^{e}(h)=u,v^{s}(h)=v}x_{h}$ $\forall u,v\in\mathcal{R}:u<v,$ (9)
@@
The objective function (7) maximises the total weight, constraints (8) enforce that all RDPs are involved in at most one cycle and constraints (9) ensure that every selected half-cycle is matched by another compatible half-cycle.
@@
Example 1.
(continued) Consider the instance presented in Figure 3 for $K=4$. If we do not change the ordering of the vertices, we have =$\mathcal{H}=\{\left\langle 1,4\right\rangle,\left\langle 4,1\right\rangle,\left\langle 4,2\right\rangle,\left\langle 1,2,3\right\rangle,\left\langle 2,3,4\right\rangle,\left\langle 3,4,1\right\rangle\}$ and $\mathcal{H}_{1}=\emptyset$. Therefore, HCF-CYCLE contains $6$ variables, one for each half-cycle. For instance, cycle $\left\langle 1,2,3,4,1\right\rangle$ is obtained by setting $x_{\left\langle 1,2,3\right\rangle}=x_{\left\langle 3,4,1\right\rangle}=1$. We refer to Delorme et al., (2023) for an example that illustrates that HCF-CYCLE typically has fewer variables than CF-CYCLE for $K\geq 4$.
@@
We consider a set $\mathcal{H}^{\prime}_{\mathcal{N}}$ of first half-chains and a set $\mathcal{H}^{\prime}_{\tau}$ of second half-chains that together constitute the set of all half-chains $\mathcal{H}^{\prime}=\mathcal{H}^{\prime}_{\mathcal{N}}\cup\mathcal{H}^{\prime}_{\tau}$. A first half-chain $h_{1}$ of length $\ell_{1}$ consists of an NDD $u_{1}\in\mathcal{N}$ followed by $\ell_{1}$ RDPs $u_{2},\ldots,u_{\ell_{1}+1}\in\mathcal{R}$, whereas a second half-chain $h_{2}$ of length $\ell_{2}$, consists of $\ell_{2}$ RDPs $v_{1},\ldots,v_{\ell_{2}}\in\mathcal{R}$ followed by the terminal node $\tau$. Reusing the notation $v^{s}(h)$, $v^{e}(h)$ and $\mathcal{V}^{m}(h)$, a first half-chain $h_{1}\in\mathcal{H}^{\prime}_{\mathcal{N}}$ of length $\ell_{1}$ and a second half-chain $h_{2}\in\mathcal{H}^{\prime}_{\tau}$ of length $\ell_{2}$ are compatible if (i) $v^{e}(h_{1})=v^{s}(h_{2})$, (ii) $\mathcal{V}^{m}(h_{1})\cap\mathcal{V}^{m}(h_{2})=\emptyset$, and (iii) $\ell_{1}+\ell_{2}\leq L$. The weight of a half-chain $h$ is again denoted by $\omega_{h}$. To reduce symmetry, we require that the lengths of the first and second half-chains satisfy $\ell_{1}\leq\left\lfloor L/2\right\rfloor$ and $\ell_{2}\leq\left\lceil L/2\right\rceil$, respectively. In addition, we only consider a first half-chain $h_{1}\in\mathcal{H}^{\prime}_{\mathcal{N}}$ of length $\ell$ if there exists at least one compatible second half-chain $h_{2}\in\mathcal{H}^{\prime}_{\tau}$ of length $\ell$ or $\ell+1$, and conversely we only consider a second half-chain $h_{2}\in\mathcal{H}^{\prime}_{\tau}$ of length $\ell$ if there exists at least one compatible first half-chain $h_{1}\in\mathcal{H}^{\prime}_{\mathcal{N}}$ of length $\ell$ or $\ell-1$.
Analogously to HCF-CYCLE, we introduce a binary decision variable $y_{h}$ for every half-chain $h\in\mathcal{H}^{\prime}$, taking value $1$ if half-chain $h$ is selected, and value 00 otherwise. HCF-CHAIN can then be defined as:
[…]
@@
The objective function (11) maximises the total weight, constraints (12) and (13) enforce that all NDDs and RDPs, respectively, are involved in at most one chain, and constraints (14) ensure that every selected first half-chain is matched by a compatible second half-chain.
If chains of length $1$ are considered, we add to $\mathcal{H}^{\prime}$ the set $\mathcal{H}^{\prime}_{\mathcal{N}\tau}$ consisting of the chains $\left\langle v,\tau\right\rangle$ for all $v\in\mathcal{N}$, and we introduce an additional binary decision variable $y_{h}$ for every $h\in\mathcal{H}^{\prime}_{\mathcal{N}\tau}$. Moreover, we replace the set $\mathcal{H}^{\prime}_{\mathcal{N}}$ under the summation sign appearing in constraints (12) by the set $\mathcal{H}^{\prime}_{\mathcal{N}}\cup\mathcal{H}^{\prime}_{\mathcal{N}\tau}$. We also present an alternative method in Appendix A.6.
Example 1.
(continued) Consider the instance presented in Figure 3 for $L=4$. We have $\mathcal{H}^{\prime}_{\mathcal{N}}=\{\left\langle 5,1\right\rangle,\left\langle 5,4\right\rangle,\left\langle 6,2\right\rangle,\left\langle 5,1,2\right\rangle,\left\langle 5,1,4\right\rangle,\left\langle 6,2,3\right\rangle\}$, $\mathcal{H}^{\prime}_{\tau}=\{\left\langle 1,\tau\right\rangle,\left\langle 2,\tau\right\rangle,\left\langle 4,\tau\right\rangle,\left\langle 1,2,\tau\right\rangle,\left\langle 1,4,\tau\right\rangle,\newline \left\langle 2,3,\tau\right\rangle,\left\langle 3,4,\tau\right\rangle,\left\langle 4,1,\tau\right\rangle,\left\langle 4,2,\tau\right\rangle\}$ and $\mathcal{H}^{\prime}_{\mathcal{N}\tau}=\{\left\langle 5,\tau\right\rangle,\left\langle 6,\tau\right\rangle\}$. Therefore, HCF-CHAIN contains $17$ variables, one for each half-chain and chain of length $1$. For instance, chain $\left\langle 5,1,2,3,\tau\right\rangle$ is obtained by setting $y_{\left\langle 5,1,2\right\rangle}=y_{\left\langle 2,3,\tau\right\rangle}=1$.

### v2

@@ 3.2 Introduction to ILP Formulations for KE-Opt
As chains were only introduced in practice later, the version of KE-Opt with chains has received less attention. To address this gap, we also show in Sections 3.3-3.7 how each cycle model can be adapted to a chain model that models a set of cardinality-constrained vertex-disjoint chains, instead. The resulting models are denoted as CF-CHAIN, HCF-CHAIN, EF-CHAIN, EEF-CHAIN and PIEF-CHAIN. For EF-CHAIN and EEF-CHAIN we consider two variants each, which are distinguished by adding suffix -EXP or -MTZ. Throughout the following sections, we assume that $L$ is finite, and we discuss in Appendix H how each of these models can be adapted to the case where $L=\infty$ (meaning that the maximum chain length is unbounded).
Even though there are no KEPs in practice that allow only for chains, separately studying cycle models and chain models enables a comprehensive view of modelling cycles and chains by merging any cycle model with any chain model to obtain a combined model that incorporates a set of cardinality-constrained vertex-disjoint cycles and chains. This works as follows: simply keep the variables and constraints of both the chosen cycle model and the chain model, sum the objective functions, and for each RDP $v\in\mathcal{R}$, merge the constraint in the cycle model stating that $v$ can be involved in at most one cycle with the corresponding constraint in the chain model stating that $v$ can be involved in at most one chain to obtain a constraint ensuring that $v$ can be involved in at most one cycle or chain.
@@
Cycle models Chain models
Model Section(s) Literature Model Section(s) Literature
CF-CYCLE 3.3 Abraham et al., (2007) / Roth et al., (2007) CF-CHAIN 3.3 Constantino et al., (2013)
HCF-CYCLE 3.4 Delorme et al., (2023) HCF-CHAIN* 3.4 Delorme et al., (2023)
EF-CYCLE 3.5 Abraham et al., (2007) / Roth et al., (2007) EF-CHAIN-EXP* 3.5 Constantino et al., (2013) & Anderson et al., 2015b
@@
EEF-CYCLE 3.6 Constantino et al., (2013) EEF-CHAIN-EXP 3.6/A Anderson et al., 2015b
EEF-CHAIN-MTZ* 3.6/A Mak-Hau, (2017)
PIEF-CYCLE 3.7 Dickerson et al., (2016) PIEF-CHAIN 3.7 Dickerson et al., (2016)
Hybrid model: EF-HYBRID*, Section(s): 3.5/C, Literature: Constantino et al., (2013) & Anderson et al., 2015b
@@ 3.3 Cycle Formulation
Let $\mathcal{C}_{\leq K}$ be the set of all cycles $c$ in $\mathcal{G}$ of length at most $K$. For every cycle $c\in\mathcal{C}_{\leq K}$ we assume that the first vertex is the lowest-indexed one (to avoid repetitions due to symmetry). Moreover, let $\mathcal{V}(c)$ and $\mathcal{A}(c)$ denote the sets of vertices and arcs in $c$, and let $\omega_{c}=\sum_{(u,v)\in\mathcal{A}(c)}w_{uv}$ denote the total weight of $c$. We introduce a binary decision variable $x_{c}$ for every $c\in\mathcal{C}_{\leq K}$, taking value $1$ if cycle $c$ is selected, and value 0 otherwise. CF-CYCLE can then be defined as follows:
$(\texttt{CF-CYCLE})\quad$ $\max$ $\sum_{c\in\mathcal{C}_{\leq K}}\omega_{c}x_{c}$ (1)
@@
The objective function (1) maximises the total weight and constraints (2) enforce that each RDP is involved in at most one cycle.
The adaptation of CF-CYCLE to model chains instead of cycles is straightforward, and was first described in a general setting by Constantino et al., (2013). Defining $\mathcal{C}^{\prime}_{\leq L}$ as the set of all chains $c$ in $\mathcal{G}$ of length at most $L$, and defining a binary variable $y_{c}$ for every $c\in\mathcal{C}^{\prime}_{\leq L}$, taking value $1$ if chain $c$ is selected and value 0 otherwise, the following model is obtained:
$(\texttt{CF-CHAIN})\qquad$ $\max$ $\sum_{c\in\mathcal{C}^{\prime}_{\leq L}}\omega_{c}y_{c}$ (4)
@@
In Appendix G we explain that the variables associated to chains of length $1$ can be omitted.
@@
The simplest approach is to enumerate all cycles and/or chains, as did, for instance, Roth et al., (2007), Manlove and O’Malley, (2014), Constantino et al., (2013), Anderson et al., 2015b and Dickerson et al., (2016). However, this approach is only viable when $K$ and $L$ are sufficiently small and the compatibility graph is sufficiently small and sparse. For example, Constantino et al., (2013) showed that CF-CYCLE could solve some low-density instances with $|\mathcal{R}|=1000$, $K=3$ and $L=0$, whereas the number of cycles was too high for high-density instances with $|\mathcal{R}|=50$, $K\geq 4$ and $L=0$. These papers deal with chains in different ways. Roth et al., (2007) considered the cycles-only case. Manlove and O’Malley, (2014) studied the case where $K=L$, and applied the chain-to-cycle transformation, after which CF-CYCLE was used. Constantino et al., (2013) extended this to the case where $K\neq L$, which essentially comes down to using the combined model CF-CYCLE + CF-CHAIN. Finally, Anderson et al., 2015b and Dickerson et al., (2016) combined CF-CYCLE with a different chain model, namely a variant of EF-CHAIN-EXP (see Section 3.5) and PIEF-CHAIN (see Section 3.7), respectively. This was motivated by the fact that the number of variables (i.e., chains) in CF-CHAIN typically dominates the number of variables (i.e., cycles) in CF-CYCLE.
@@
Several other improvements for B&P algorithms were pursued. First, different branching rules have been proposed. For example, Abraham et al., (2007) branched directly on fractional cycle/chain variables, whereas Glorie et al., 2014b branched on the underlying fractional arcs, which ensures that the depth of the branching tree is polynomially bounded. Moreover, Klimentova et al., (2014) introduced a new decomposition model for the master problem, called the Disagreggated Cycle Decomposition Model, which was improved by Riascos-Álvarez et al., (2024) using feedback vertex sets. Furthermore, Lam and Mak-Hau, (2020) and Petris et al., (2024) considered branch-and-price-and-cut algorithms, in which one progressively adds cutting planes to strengthen the LP relaxation of the restricted master problem. Whereas Lam and Mak-Hau, (2020) report that their considered valid inequalities are rather ineffective in strengthening the LP relaxation, the inequalities introduced by Petris et al., (2024) are shown to be more effective. Moreover, Riascos-Álvarez et al., (2024) introduced a new upper bound based on Lagrangian relaxation. Very recently, Arslan et al., (2024) proposed two B&P algorithms (one based on CF-CYCLE+CF-CHAIN and one based on CF-CYCLE+PIEF-CHAIN) that combine many of the improvements that were proposed over the years. As shown also by our computational experiments (see Section 4), these algorithms are the current state-of-the-art for KE-Opt on unweighted instances among all publicly available methods.
Delorme et al., (2023) and Delorme et al., (2024) were the first to apply RCVF to KE-Opt. Less involved than B&P, RCVF is a general technique that can be applied to any ILP model, which works especially well when the model’s LP relaxation bound is close to the optimal integer solution value (as is the case for CF-CYCLE and CF-CHAIN, see Section 3.8). The idea is to deactivate variables for which, based on their reduced cost, it can be deduced that their value will never be $1$ or higher in a solution whose value is equal to some upper bound. This upper bound is initially set to the value of the LP relaxation rounded down, and then progressively decreased within a destructive bound framework until an optimal solution is found.
@@ 3.4 Half-Cycle Formulation
Formally, a half-cycle $h$ of length $k$, denoted as $h=\left\langle v_{1},v_{2},\ldots,v_{k+1}\right\rangle$, is a simple path in $\mathcal{G}$ consisting of $k+1$ RDPs $v_{1},\ldots,v_{k+1}\in\mathcal{R}$ such that there is an arc $(v_{i},v_{i+1})\in\mathcal{A}_{\mathcal{R}}$ for every $i=1,\ldots,k$. Let $v^{s}(h)=v_{1}$ and $v^{e}(h)=v_{k+1}$ denote the start and end vertex of half-cycle $h$, respectively, and let $\mathcal{V}^{m}(h)=\{v_{2},\ldots,v_{k}\}$ denote the set of intermediate vertices (when $k\geq 2$). Moreover, letting $\mathcal{A}(h)$ denote the set of arcs in half-cycle $h$, its total weight is given by $\omega_{h}=\sum_{(i,j)\in\mathcal{A}(h)}w_{ij}$. Using this notation, we have that two half-cycles, say $h_{1}$ of length $k_{1}$ and $h_{2}$ of length $k_{2}$, are compatible (i.e., they may be combined to a full cycle of weight $\omega_{h_{1}}+\omega_{h_{2}}$) if $v^{e}(h_{1})=v^{s}(h_{2})$, $v^{e}(h_{2})=v^{s}(h_{1})$, $\mathcal{V}^{m}(h_{1})\cap\mathcal{V}^{m}(h_{2})=\emptyset$, and $k_{1}+k_{2}\leq K$.
Let $\mathcal{H}$ denote the set of all half-cycles. Given an ordering of the vertices in $\mathcal{V}$, it is sufficient to only consider half-cycles $h$ of length $k$ that satisfy the following conditions: (i) $k\leq\left\lceil K/2\right\rceil$, (ii) either $v^{s}(h)$ or $v^{e}(h)$ is the smallest indexed vertex among vertices in $h$, but it must be $v^{s}(h)$ if $K$ is odd and $k=\left\lceil K/2\right\rceil$, and (iii) there exists at least one compatible half-cycle of length $k$ or k1k-1 if $v^{s}(h)<v^{e}(h)$, or length $k$ or $k+1$ if $v^{s}(h)>v^{e}(h)$ [^8] In our implementation, we first construct all half-cycles satisfying conditions (i) and (ii), after which we remove all half-cycles that do not satisfy a relaxed version of condition (iii). Namely, we ignore the requirement that two compatible half-cycles $h_{1}$ and $h_{2}$ must satisfy $\mathcal{V}^{m}(h_{1})\cap\mathcal{V}^{m}(h_{2})=\emptyset$.. There are multiple ways to determine an ordering of the vertices, but Delorme et al., (2023) proposed to sort the vertices in descending order of total vertex degree, as this typically leads to a smaller number of half-cycles.
Introducing a binary decision variable $x_{h}$ for every half-cycle $h\in\mathcal{H}$, taking value $1$ if half-cycle $h$ is selected, and value 0 otherwise, HCF-CYCLE can be defined as the following ILP:
$(\texttt{HCF-CYCLE})\quad$ $\max$ $\sum_{h\in\mathcal{H}}\omega_{h}x_{h}$ (7)
s.t. $\sum_{h\in\mathcal{H}:v\in\mathcal{V}^{m}(h)\cup\{v^{s}(h)\}}\hskip-3.0ptx_{h}\leq 1$ $\forall v\in\mathcal{R},$ (8)
$\sum_{h\in\mathcal{H}:v^{s}(h)=u,v^{e}(h)=v}\hskip-3.0ptx_{h}=\sum_{h\in\mathcal{H}:v^{e}(h)=u,v^{s}(h)=v}\hskip-3.0ptx_{h}$ $\forall u,v\in\mathcal{R}:u<v,$ (9)
@@
The objective function (7) maximises the total weight, constraints (8) enforce that each RDP is involved in at most one cycle and constraints (9) ensure that every selected half-cycle is matched by another compatible half-cycle.
@@
We consider a set $\mathcal{H}^{\prime}_{\mathcal{N}}$ of first half-chains and a set $\mathcal{H}^{\prime}_{\tau}$ of second half-chains that together constitute the set of all half-chains $\mathcal{H}^{\prime}=\mathcal{H}^{\prime}_{\mathcal{N}}\cup\mathcal{H}^{\prime}_{\tau}$. A first half-chain $h_{1}$ of length $\ell_{1}$ consists of an NDD $u_{1}\in\mathcal{N}$ followed by $\ell_{1}$ distinct RDPs $u_{2},\ldots,u_{\ell_{1}+1}\in\mathcal{R}$, whereas a second half-chain $h_{2}$ of length $\ell_{2}$, consists of $\ell_{2}$ distinct RDPs $v_{1},\ldots,v_{\ell_{2}}\in\mathcal{R}$ followed by the terminal node $\tau$. Reusing the notation $v^{s}(h)$, $v^{e}(h)$ and $\mathcal{V}^{m}(h)$, a first half-chain $h_{1}\in\mathcal{H}^{\prime}_{\mathcal{N}}$ of length $\ell_{1}$ and a second half-chain $h_{2}\in\mathcal{H}^{\prime}_{\tau}$ of length $\ell_{2}$ are compatible if (i) $v^{e}(h_{1})=v^{s}(h_{2})$, (ii) $\mathcal{V}^{m}(h_{1})\cap\mathcal{V}^{m}(h_{2})=\emptyset$, and (iii) $\ell_{1}+\ell_{2}\leq L$. The weight of a half-chain $h$ is again denoted by $\omega_{h}$. To reduce symmetry, we require that the lengths of the first and second half-chains satisfy $\ell_{1}\leq\left\lfloor L/2\right\rfloor$ and $\ell_{2}\leq\left\lceil L/2\right\rceil$, respectively. In addition, we only consider a first half-chain $h_{1}\in\mathcal{H}^{\prime}_{\mathcal{N}}$ of length $\ell$ if there exists at least one compatible second half-chain $h_{2}\in\mathcal{H}^{\prime}_{\tau}$ of length $\ell$ or $\ell+1$, and conversely we only consider a second half-chain $h_{2}\in\mathcal{H}^{\prime}_{\tau}$ of length $\ell$ if there exists at least one compatible first half-chain $h_{1}\in\mathcal{H}^{\prime}_{\mathcal{N}}$ of length $\ell$ or $\ell-1$.
Analogously to HCF-CYCLE, we introduce a binary decision variable $y_{h}$ for every half-chain $h\in\mathcal{H}^{\prime}$, taking value $1$ if half-chain $h$ is selected, and value 0 otherwise. HCF-CHAIN can then be defined as:
$(\texttt{HCF-CHAIN})\quad$ $\max$ $\sum_{h\in\mathcal{H}^{\prime}}\omega_{h}y_{h}$ (11)
@@
The objective function (11) maximises the total weight, constraints (12) and (13) enforce , respectively, that each NDD and each RDP is involved in at most one chain, and constraints (14) ensure that every selected first half-chain is matched by a compatible second half-chain.
If chains of length $1$ are considered, we add to $\mathcal{H}^{\prime}$ the set $\mathcal{H}^{\prime}_{\mathcal{N}\tau}$ consisting of the chains $\left\langle v,\tau\right\rangle$ for all $v\in\mathcal{N}$, and we introduce an additional binary decision variable $y_{h}$ for every $h\in\mathcal{H}^{\prime}_{\mathcal{N}\tau}$. Moreover, we replace the set $\mathcal{H}^{\prime}_{\mathcal{N}}$ under the summation sign appearing in constraints (12) by the set $\mathcal{H}^{\prime}_{\mathcal{N}}\cup\mathcal{H}^{\prime}_{\mathcal{N}\tau}$. We also present an alternative method in Appendix G.